Grouper Corporation purchased machinery on January 1, 2022, at a cost of $264,000. The estimated useful life of the machineryPrepare separate depreciation schedules for the machinery using the straight-line method, and the declining-balance method usPrepare separate depreciation schedules for the machinery using the straight-line method, and the declining-balance method usDOUBLE-DECLINING-BALANCE DEPRECIATION Computation Years Boo * Depreciation Rate = Annual Depreciation Expense Accumulate DeprDOUBLE-DECLINING-BALANCE DEPRECIATION utation End of Year * Depreciation Rate = Annual Depreciation Expense Accumulated Depre
Answer #1

a)Depreciable cost= cost -salvage value

               = 264000-31400

               = 232600

Depreciation rate= 1/useful life

            = 1/4

             = .25 or 25%

STRAIGHT LINE DEPRECIATION
COMPUTATION END OF YEAR
Depreciable cost * Depreciation rate = Depreciation expense Accumulated depreciation Book value at end
2022 232600 25% 58150 58150 264000-58150= 205850
2023 232600 25% 58150 58150+58150= 116300 264000-116300= 147700
2024 232600 25% 58150 116300+58150= 174450 264000-174450= 89550
2025 232600 25% 58150 174450+58150= 232600 264000-232600= 31400
232600

b)

Depreciation rate= 2/useful life

            = 2/4

             = .50 or 50%

DOUBLE DECLINING DEPRECIATION
COMPUTATION END OF YEAR
Book value beginning of year * Depreciation rate = Depreciation expense Accumulated depreciation Book value at end
2022 264000 50% 132000 132000 264000-132000=132000
2023 132000 50% 66000 132000+66000=198000 264000-198000= 66000
2024 66000 50% 33000 198000+33000= 231000 264000-231000=33000
2025 33000 1600/33000=4.85% 33000-31400=1600 231000+1600=232600 31400
232600

In year 2025,Charging of actual depreciation of 33000*50% = 16500 will make book value at end equals to 16500 which is below minimum salvage value required .Thus depreciation for the year will be restricted such that book value at end is equals to salvage value of 31400